OpenStudy (anonymous):

convert from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places ,as needed. 36°

OpenStudy (anonymous):

The answer is π/5. I want to know how this is the answer.

OpenStudy (lgbasallote):

multiply by pi/180 \[\huge 36^o \times \frac{\pi}{180^o } \implies \cancel{36^o} \times \frac{\pi}{\cancel{180^o }_5} \implies \frac \pi 5\] does that help?

OpenStudy (lgbasallote):

btw...if you're wondering...when you want to convert degrees into radians..you ALWAYS multiply the given degree by pi/180